OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 2. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (anonymous):

Hint: Do you know how to remove the base-16 logarithm from the left-hand-side of the equation?

OpenStudy (anonymous):

32 = 16^(x + 2) Note that 32 = 2^5 and 16 = 2^4. Then, 2^5 = 2^(4(x + 2)) 2^5 = 2^(4x + 8) Set the exponents = to each other and solve for x: 5 = 4x + 8 4x = 5 - 8 4x = -3 x = -3/4